OpenStudy (anonymous):

Convert 75mph to km/s. According to the text, the answer is 0.034 km/s (or 3.4x10^-2 km/s). I have no idea how you arrive at the answer. I know you have to keep canceling out the units until you've reached the desired one, but I can't figure out how. 1. 75 mi/hr 2. Convert miles to feet and cancel out miles >> 75 mi/hr x 5280 ft/mi But where did the 5280 come from?

OpenStudy (phi):

mph is short for miles per hour, or , written as a fraction \[ \frac{miles}{hour} \] you want km/sec tackle the top and bottom separately first the top: miles you want km the trick is to think like in algebra (if you know how) \[ miles \cdot \frac{km}{miles} \] and like in algebra, the miles can be cancelled \[\cancel{ miles} \cdot \frac{km}{\cancel{ miles}} =km\] that means we need a fraction that tells us the number of km in a mile we look that up (or remember it) \[ \frac{1.609344 \ km}{1 \ mile} \] use that fraction to convert your miles: \[ 75 \ miles \cdot \frac{1.609344 \ km}{1 \ mile} \] can you find how many km that is ?

OpenStudy (phi):

once you can do this, we can do the same thing using feet. same idea, but more steps